### v2

# Improving Vision Transformers by Overlapping Heads in Multi-Head Self-Attention

Abstract
Vision Transformers have made remarkable progress in recent years, achieving state-of-the-art performance in most vision tasks. A key component of this success is due to the introduction of the Multi-Head Self-Attention (MHSA) module, which enables each head to learn different representations by applying the attention mechanism independently. In this paper, we empirically demonstrate that Vision Transformers can be further enhanced by overlapping the heads in MHSA. We introduce Multi-Overlapped-Head Self-Attention (MOHSA), where heads are overlapped with their two adjacent heads for queries, keys, and values, while zero-padding is employed for the first and last heads, which have only one neighboring head. Various paradigms for overlapping ratios are proposed to fully investigate the optimal performance of our approach. The proposed approach is evaluated using five Transformer models on four benchmark datasets and yields a significant performance boost. The source code will be made publicly available upon publication.
Index Terms:
Vision Transformers, Attention Mechanism, Multi-Head Self-Attention
I Introduction
Since the introduction of Transformer [1] from language to vision, Vision Transformers have gradually dominated many vision tasks such as image recognition [2][3][4] and object detection [5][6][7][8]. Due to the global attention mechanism in Transformers, Vision Transformers could yield better performance in vision tasks when the training data is abundant and the training epochs are enough. Although the convergence rate of Vision Transformers is frequently slower than their Convolutional counterparts, global information communication is one of the key elements for the success of Transformer models.
Multi-Head Self-Attention (MHSA) [1] is the core module for Transformer models so that the self-attention mechanism can be implemented in multiple heads to learn different representations. The self-attention mechanism is originally designed for processing language data to capture the long-range relationship between words in language. When self-attention is applied to vision tasks for image classification and recognition, each image patch token could interact with all other patch tokens for the global views of the image, which is significant for the excellent performance of Vision Transformers. MHSA represents that the queries, keys and values are split into different heads and the self-attention is computed in each head independently. Transformer models with Multi-Head Attention frequently yield better performance than those with Single-Head Attention. Thus most Transformer models utilize Self-Attention with multiple heads for better representation learning and performance. Even though MHSA boosts the performance of the Transformer models, the relationships and interactions between different heads are rarely investigated.
The success of Transformer models mainly lies on the effective information exchange between different tokens so that each token can have global views of the context information. Due to the superior performance of MHSA, most Transformer models utilize MHSA by default. However, the queries, keys and values are divided for each head without overlapping and there is no information exchange when the attention is computed in each head. In other words, when calculating the attention in the current head, it does not have the information in other heads. Although the tokens will be processed by linear projections after the attention, the information exchange is only limited to each token.
Figure 1: The proposed multi-overlapped-head method (blue) vs the original multi-head method (green). Instead of hard division of the heads, our approach softly splits the heads by overlapping each head with its neighboring heads.
In this paper, we claim that information exchange during attention calculation in each head can improve the performance of Vision Transformers. This can be realized by overlapping queries, keys and values in each head with queries, keys and values of neighboring heads. For this purpose, we propose Multi-Overlapped-Head Self-Attention (MOHSA) to improve the Multi-Head Self-Attention mechanism by overlapping the heads so that $Q$, $K$, and $V$ in each head are overlapped by $Q$, $K$, and $V$ of their neighboring heads when the attention is calculated, as illustrated in Fig. 1. By overlapping the heads, the information in other heads could also be involved in the calculation of the attention in the current head. Since the overlapping would slightly increase the dimension of the tokens after concatenating tokens from different heads, the linear projections would decrease the dimension of the tokens to the original size. The information communication between the heads could yield better performance for Vision Transformers. Moreover, we design various paradigms for overlapping ratios to investigate the best performance of our proposed approach. We explore several Vision Transformer models on CIFAR-10 [9], CIFAR-100 [9], Tiny-ImageNet [10] and ImageNet-1k [11] to verify the effectiveness of our proposed MOHSA.
Our major contributions are summarized below.
  • We propose Multi-Overlapped-Head Self-Attention (MOHSA) and demonstrate that Vision Transformer models could be improved by overlapping the queries, keys and values of the current head with the queries, keys and values of adjacent heads when the attention is computed. Several Vision Transformer models are exploited on various datasets to demonstrate the effectiveness of our proposed method.
  • Various variants based on the overlap dimensions are proposed to thoroughly investigate the optimal performance of MOHSA. The proposed MOHSA could be integrated into Vision Transformer models to enhance their performance with negligible overhead.
  • To the best of our knowledge, our work is the first to investigate the overlapping approach between different heads for MHSA when the attention is calculated.
II Related Work
II-A Vision Transformers
Vision Transformer [2] utilizes the image patches which are embedded as tokens and a class token as the token inputs for the Transformer encoder to recognize the images. Swin-Transformer [3] reduces the computational cost by limiting the implementation of attention calculation in each window and expands the view of the tokens by shifting the windows. Shuffle Transformer [12] proposes spatial shuffle to exchange information across windows. MSG-Transformer [13] harnesses MSG tokens to represent local windows and the information communication between windows is implemented with MSG tokens. PVT [14] designs a Vision Transformer with a hierarchical structure without convolutions by shrinking the feature maps gradually. CaiT [15] and DeepViT [16] investigate Vision Transformers with deeper layers. A multi-layer dense attention decoder was proposed in [17] for segmentation. MobileViT series [18][19][20] and EdgeViTs [21] explore Vision Transformers on mobile-level applications. Some works [22][23][24][25][26] introduce convolutions into the Vision Transformers to take advantage of both convolutions and Transformers. PiT [27] explores the possibility of using individual pixels instead of patches as the tokens for Vision Transformers. ViTAR [28] adjusts Vision Transformers on various image resolutions. PartialFormer [29] proposes partial attention for more efficient Vision Transformers on various vision tasks.
II-B Attention Heads Interaction
DeepViT [16] explores a deeper layer Vision Transformer by proposing re-attention that mixes the attention maps among all heads with a learnable matrix before multiplying with the values. The re-attention is similar to talking-heads attention [30] which is originally employed for language tasks and also utilized by CaiT [15]. Talking-heads attention [30] applies learnable linear projections to the heads in Multi-Head Attention before and after the softmax function to exchange the information for the attention maps between heads in the attention module.
The aforementioned methods mix attention maps using linear projections, allowing information exchange between heads only after the attention maps are calculated. However, the mixed attention maps are then applied to values that do not have information exchange between heads. In contrast, we propose a method that enables communication between heads during the attention computation by overlapping the heads with $Q$, $K$, and $V$. Although this approach introduces a slight increase in computation and parameters, it significantly enhances the performance of Vision Transformer models across various datasets.
Figure 2: The Transformer encoder for a typical Transformer model. The Transformer encoder is exploited in Vision Transformer for image classification. $N$ indicates the number of layers for the Transformer encoder.
III Approach
Figure 3: Our proposed Multi-Overlapped-Head Self-Attention. MHSA represents the original implementation of Multi-Head Self-Attention with hard division of heads and MOHSA indicates our proposed Multi-Overlapped-Head Self-Attention with soft division of heads. In the original Vision Transformer (left), $Q$, $K$, and $V$ are split for different heads and the attention is computed for each head independently. To exchange the information between heads when the attention is calculated, we propose to overlap $Q$, $K$, $V$ with $Q$, $K$, and $V$ in adjacent heads (right). Since overlapped heads would slightly increase the number of dimensions, the projection matrix would project the concatenated heads to the original token dimension.
III-A Multi-Head Self-Attention
Transformer [1] has an attention mechanism to compute the long-range relationship between tokens. The attention calculation includes queries, keys, values and the dimension of queries and keys $d_{k}$. The queries and keys are implemented dot product to calculate the weights which are utilized to the values to compute the final results. The matrix format of attention calculation [1] is illustrated in Eq. (1). $\sqrt{d_{k}}$ is employed to prevent the large value input for the softmax function after computing the dot product [1].
$Attention(Q,K,V)=Softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
Multi-Head Self-Attention [1] is utilized in the Transformer model for better performance so that different heads can learn different representations, which is better than one-head attention [1]. The attention in each head is illustrated in Eq. (2). $Q$, $K$, and $V$ are divided as $Q_{i}$, $K_{i}$, and $V_{i}$ in each head respectively and the attention calculation is implemented in each head independently for different aspects of learning.
$Head_{i}=Attention(Q_{i},K_{i},V_{i})$ (2)
$MHSA(Q,K,V)=Concate(Head_{1},...,Head_{n})W$
Finally, MHSA could be represented as Eq. (3). The results of all heads are concatenated and $W$ is the projection matrix.
Vision Transformer [2] is the Transformer model applied to vision tasks, mainly image classification and recognition. For Vision Transformer, only the encoder is utilized for feature extraction, as demonstrated in Fig. 2. The images are divided as same-size patches which are embedded as tokens for the Vision Transformer. Layer Normalization [31] is often employed as the normalization before MHSA and FFN. Vision Transformer [2] could be applied to vision tasks and achieve significant performance on vision tasks due to the effective information exchange for attention mechanism. the self-attention mechanism calculates the dot product of each token with all other tokens effectively exchanging the information between them. In vision tasks like image classification and recognition, each patch token has a global view of the image, which is important for each patch to attain the context information.
Nonetheless, the attention is implemented independently in each head and one head does not have the attention information of other heads when the attention is computed. Although the projection matrix is implemented after all the heads are concatenated, only the information in each token is exchanged. There is no information exchange when computing the attention in each head. Thus we propose to overlap the information of neighboring heads to enhance the information exchange when the attention is computed in each head.
III-B Multi-Overlapped-Head Self-Attention
Based on the aforementioned analysis, we have proposed a simple yet effective approach that improves Multi-Head Self-Attention to enhance the performance of Vision Transformer. To make information exchange between heads, we exploit soft division instead of hard division when $Q$, $K$, and $V$ are divided into different heads. The process could be illustrated in Eq. (4) for $Q$, $K$, and $V$ respectively. We utilize “part” to illustrate partial overlapping with adjacent heads in Eq. (4).
$\begin{split}&Q_{i}^{\prime}=Concate(part(Q_{i-1}),Q_{i},part(Q_{i+1}))\\ &K_{i}^{\prime}=Concate(part(K_{i-1}),K_{i},part(K_{i+1}))\\ &V_{i}^{\prime}=Concate(part(V_{i-1}),V_{i},part(V_{i+1}))\end{split}$ (4)
In Eq. (4), $Q_{i}$, $K_{i}$ and $V_{i}$ are the original hard division results for $Head_{i}$, which is shown in Eq. (2). For soft division, $Q_{i}^{\prime}$, $K_{i}^{\prime}$ and $V_{i}^{\prime}$ also include partial information in their neighboring heads so that $Q$, $K$, $V$ are overlapped with two adjacent heads, which is demonstrated in Fig. 3. The left figure indicates the original implementation of the hard division of $Q$, $K$, $V$ to different heads and the right figure represents our proposed implementation of the soft division of $Q^{\prime}$, $K^{\prime}$, $V^{\prime}$ to different heads. The $Q_{i}$, $K_{i}$, and $V_{i}$ would overlap with two adjacent heads to construct $Q_{i}^{\prime}$, $K_{i}^{\prime}$, and $V_{i}^{\prime}$ for calculating the attention. For the first head and the last head which only have one adjacent head, zero padding is utilized to construct two neighboring heads for the first and the last head.
$Head_{i}^{\prime}=Attention(Q_{i}^{\prime},K_{i}^{\prime},V_{i}^{\prime})$ (5)
$MOHSA(Q,K,V)=Concate(Head_{1}^{\prime},...,Head_{n}^{\prime})W^{\prime}$
After the attention for the overlapped $Q^{\prime}$, $K^{\prime}$, and $V^{\prime}$ is calculated for each head, the results are concatenated together, as illustrated in Eq. 5-6. The overlapped heads would slightly increase the dimension of the tokens after concatenation. Thus the projection matrix $W^{\prime}$ would project the concatenated dimension ($h*dim_{h}^{\prime}$) to the original token dimension ($dim$) so that the token could be fed into the next layer with the same dimension. The projection matrix $W$ in Eq. (3) projects the concatenated non-overlapped heads ($h*dim_{h}$) to the original token dimension ($dim$). Since the dimension of overlapped heads $dim_{h}^{\prime}$ is slightly larger than the dimension of non-overlapped heads $dim_{h}$ and the number of heads $h$ is unchanged, the projection matrix $W^{\prime}$ in our proposed MOHSA would have slightly more parameters than the projection matrix $W$ in the original MHSA.
TABLE I: The variants for overlapping ratios
Methods Description
inc (x layers) increase overlap dim by 1 every x layers
dec (x layers) decrease overlap dim by 1 every x layers
0-indexed start (end) overlap dim from (to) 0
1-indexed start (end) overlap dim from (to) 1
Figure 4: The illustration of the overlap dimensions. The blue parts demonstrate the original non-overlapping heads and the red parts indicate the overlapped parts from adjacent heads. The number of overlap dimensions is the overlap dimension of one side adjacent head.
III-C Overlapping Ratios
The overlapping ratios for the overlapped heads are crucial to the effectiveness of the proposed approach. In the experiments, we design several paradigms for the overlapping ratios. In this work, we utilize overlap dimensions to demonstrate the overlapping ratios, as illustrated in Fig. 4. From Fig. 4, the number of overlap dimensions for each head is the overlap dimension of its one-side adjacent head. In Fig. 4, the blue sections are the original parts of the heads and the red sections are overlapped parts with two adjacent heads. For the first and the last head which have only one side adjacent head, zero padding is exploited for the missing neighboring heads.
In addition to the fixed overlap dimensions for all layers, we also implement some variants of the overlapping ratios by changing the overlap dimensions according to the depths of the layers. In this paper, one layer includes the attention module and the FFN module, as illustrated in Fig. 2 which represents one layer of the Transformer encoder. The variants for the overlapping ratios are demonstrated in Table I. In Table I, “inc (x layers)” represents the overlap dimension is increased by 1 every x layers, and “dec (x layers)” indicates the overlap dimension is decreased by 1 every x layers, which is a reverse process of ”inc (x layers)” based on the depths of the layers. In addition, “0-indexed” illustrates the overlap dimension starts from 0 for “inc” and ends at 0 for “dec”, and “1-indexed” demonstrates the overlap dimension starts from 1 for “inc” and ends at 1 for “dec”. For instance, inc (2 layers) with “0-indexed” (represented by “inc-0 (2)” in the experiments) for a total of 12 layers has overlap dimensions (0, 0, 1, 1, 2, 2, 3, 3, 4, 4, 5, 5) for layer from 1 to 12, and dec (1 layer) with “1-indexed” (represented by “dec-1 (1)” in the experiments) for total 12 layers has overlap dimensions (12, 11, 10, 9, 8, 7, 6, 5, 4, 3, 2, 1) for layer from 1 to 12.
TABLE II: The parameter settings for the models
Models heads depths MLP ratio
ViT-Tiny 12 12 4
ViT-Small 12 12 4
CaiT-xxs12 4 12 (2) 4
CaiT-xxs24 4 24 (2) 4
Swin-Tiny (3, 6, 12, 24) (2, 2, 6, 2) 4
IV Experiments
In the experiments, we select three representative models to investigate the proposed approach: vanilla ViT [2], deeper layer Transformer CaiT [15], and window-based hierarchical Transformer Swin-Transformer [3]. More specifically, ViT-Tiny [2], ViT-Small [2], CaiT-xxs12 [15], CaiT-xxs24 [15] and Swin-Tiny [3] are selected to explore the effectiveness of the proposed approach. The parameter settings for the models are shown in Table II. For Swin-Tiny, the number of heads and depths are varied based on the stages. CaiT-xxs12 and CaiT-xxs24 have two extra class layers. The models will be trained and evaluated on CIFAR-10 [9], CIFAR-100 [9], Tiny-ImageNet [10] and ImageNet-1k [11].
The experiments are implemented by 100 epochs with 20 epochs warmup. The optimizer is AdamW [32]. The experiments for CIFAR-10 [9], CIFAR-100 [9], Tiny-ImageNet [10] are run on 4 NVIDIA P100 GPUs with total batch size 128. The experiments for ImageNet-1k [11] are conducted on 4 NVIDIA V100 GPUs with a total batch size of 512. The initial learning rate for ViT and CaiT is 0.0005. For Swin-Tiny, the initial learning rate for Tiny-ImageNet is 0.00025, and for ImageNet is 0.0005. The images are resized to 224 for the experiments. Some other settings follow the settings in Swin-Transformer [3].
TABLE III: The ablation study on ViT for CIFAR-10
Methods ViT-Tiny ViT-Small
Accuracy Params FLOPs Accuracy Params FLOPs
original 84.31 5.5M 1.3G 86.53 21.7M 4.6G
fixed 1 84.94 5.6M 1.3G 86.98 21.8M 4.7G
fixed half 85.34 6.0M 1.5G 86.84 23.4M 5.3G
inc-0 (1) 85.54 5.8M 1.4G 87.30 22.3M 4.9G
inc-0 (3) 84.83 5.6M 1.3G 87.24 21.8M 4.7G
inc-0 (6) 85.65 5.5M 1.3G 86.74 21.7M 4.6G
inc-1 (1) 85.35 5.9M 1.5G 86.98 22.4M 4.9G
inc-1 (3) 84.31 5.7M 1.3G 86.83 21.9M 4.7G
inc-1 (4) 85.18 5.6M 1.3G 86.81 21.9M 4.7G
dec-0 (1) 85.43 5.8M 1.4G 86.67 22.3M 4.9G
dec-0 (3) 84.81 5.6M 1.3G 86.35 21.8M 4.7G
dec-1 (1) 85.25 5.9M 1.5G 86.87 22.4M 4.9G
dec-1 (3) 85.12 5.7M 1.3G 86.84 21.9M 4.7G
TABLE IV: The ablation study on CaiT-xxs12 for CIFAR-10
Methods Accuracy Params FLOPs
original 77.04 6.4M 1.3G
fixed 1 78.14 6.4M 1.3G
inc-0 (1) 78.79 6.5M 1.4G
inc-0 (2) 78.63 6.5M 1.3G
inc-0 (3) 80.04 (+3.00) 6.4M 1.3G
IV-A CIFAR-10
CIFAR-10 [9] includes 50k training images and 10k testing images. Among all 60k images, there are 10 classes and each class has 6k images. The ablation study for ViT-Tiny and ViT-Small on CIFAR-10 is demonstrated in Table III. In Table III, “fixed 1” represents the overlap dimension is 1 for all layers, and “fixed half” indicates the overlap dimension is half of the head dimension for all layers. “inc” indicates the overlap dimension is increased with the depths of the layers increasing and “dec” represents the overlap dimension is decreased with the layers going deeper. The two numbers following “inc” or “dec”, for instance, “inc-0 (3)” illustrates the overlap dimension is increased by 1 every 3 layers and it is “0-indexed”. More details about the variants of overlapping ratios can be found in Sec. III-C.
The best accuracy for ViT-Tiny is 85.65 for variant “inc-0 (6)” which increases the overlap dimension by 1 every 6 layers with “0-indexed” and the best accuracy for ViT-Small is 87.30 for variant “inc-0 (1)” which increases the overlap dimension by 1 every 1 layer with “0-indexed”. For ViT-Tiny, “fixed half” demonstrates higher accuracy than “fixed 1”. For ViT-Small, “fixed half” has lower accuracy than “fixed 1”.
Although the overlapped heads would slightly increase the computations and parameters, from the experimental results we can see that the increased numbers of parameters and FLOPs are negligible. Table IV illustrates the experimental results of CaiT-xxs12 on CIFAR-10. The best result 80.04 significantly boosts the performance of CaiT-xxs12 by 3% with negligible overhead.
TABLE V: The ablation study on ViT for CIFAR-100
Methods ViT-Tiny ViT-Small
Accuracy Params FLOPs Accuracy Params FLOPs
original 61.88 5.5M 1.3G 63.78 21.7M 4.6G
fixed 1 62.17 5.6M 1.3G 63.90 21.8M 4.7G
fixed half 62.61 6.0M 1.5G 63.76 23.5M 5.3G
inc-0 (1) 62.86 5.8M 1.4G 64.83 22.3M 4.9G
inc-0 (2) 63.01 5.7M 1.3G 64.21 22.0M 4.7G
inc-1 (1) 62.80 5.9M 1.5G 64.49 22.4M 4.9G
inc-1 (2) 62.77 5.7M 1.4G 64.24 22.1M 4.8G
dec-0 (1) 62.45 5.8M 1.4G 64.62 22.3M 4.9G
dec-1 (3) 62.25 5.7M 1.3G 64.97 22.0M 4.7G
IV-B CIFAR-100
CIFAR-100 [9] includes 50k training images and 10k testing images. There are 100 classes and each class has 600 images. Table V demonstrates the experiments of ViT-Tiny and ViT-Small on CIFAR-100. The best result for ViT-Tiny is 63.01 with “inc-0 (2)” which increases the overlap dimension by 1 every 2 layers with “0-indexed” and for ViT-small is 64.97 with “dec-1 (3)” which decreases the overlap dimension by 1 every 3 layers with “1-indexed”. Both results improve the models by more than 1% than the original models. Similar to the results on CIFAR-10, “fixed half” has higher accuracy for ViT-Tiny, and “fixed 1” illustrates better performance for ViT-Small. In addition, “fixed half” does not demonstrate effectiveness for ViT-Small.
Table VI illustrates the performance of CaiT on CIFAR-100. For CaiT-xxs12, using 1 overlap dimension for all layers could boost the accuracy by nearly 2%. For CaiT-xxs24, the accuracy is significantly enhanced by 5% with “inc-1 (3)” which increases the overlap dimension by 1 every 3 layers with “1-indexed”. Additionally, “inc-1 (3)” is much better than using half dimension of the head as the overlap dimension which utilizes more parameters and FLOPs.
TABLE VI: The ablation study on CaiT for CIFAR-100
Models Methods Accuracy Params FLOPs
CaiT-xxs12 original 50.95 6.4M 1.3G
fixed 1 52.94 (+1.99) 6.4M 1.3G
CaiT-xxs24 original 50.94 11.8M 2.5G
fixed half 52.47 12.7M 3.1G
inc-1 (1) 55.84 12.2M 2.8G
inc-1 (3) 55.94 (+5.00) 11.9M 2.6G
IV-C Tiny-ImageNet
TABLE VII: The experimental results on Tiny-ImageNet
Models Methods Accuracy Params FLOPs
ViT-Tiny original 50.79 5.6M 1.3G
fixed 1 51.44 5.6M 1.3G
fixed half 52.32 6.0M 1.5G
inc-0 (1) 51.45 5.9M 1.4G
inc-1 (2) 52.40 (+1.61) 5.7M 1.4G
dec-0 (1) 52.21 5.9M 1.4G
dec-1 (2) 52.20 5.7M 1.4G
ViT-Small original 54.53 21.7M 4.6G
fixed 1 55.15 21.8M 4.7G
fixed half 54.65 23.5M 5.3G
inc-0 (2) 55.30 22.0M 4.7G
inc-1 (1) 55.80 (+1.27) 22.4M 4.9G
dec-1 (3) 55.46 22.0M 4.7G
CaiT-xxs12 original 42.66 6.5M 1.3G
fixed 1 45.05 (+2.39) 6.5M 1.3G
CaiT-xxs24 original 42.46 11.8M 2.5G
fixed 1 46.04 11.8M 2.6G
inc-0 (2) 46.98 12.0M 2.7G
inc-1 (1) 49.87 (+7.41) 12.2M 2.8G
Swin-Tiny original 57.04 27.7M 4.5G
fixed 1 58.27 27.8M 4.5G
inc-0 (6) 58.67 27.8M 4.5G
inc-1 (3) 58.58 28.1M 4.6G
dec-0 (1) 58.59 28.0M 4.7G
dec-1 (1) 58.90 (+1.86) 28.1M 4.7G
Tiny-ImageNet [10] includes 100k training images and 10k validation images. There are 200 classes in the dataset. Each class has 500 images for training and 50 images for validation. The experimental results on Tiny-ImageNet are demonstrated in Table VII. The accuracy of ViT-Tiny could be enhanced by 1.61% utilizing “inc-1 (2)” that increases the overlap dimension by 1 every 2 layers with “1-indexed”. ViT-Small is boosted by 1.27% with “inc-1 (1)” that increases the overlap dimension by 1 every 1 layer with “1-indexed”. Moreover, using half of the head dimension as the overlap dimension is better than using 1 as the overlap dimension for all layers for ViT-Tiny, while it is not the case for ViT-Small.
For CaiT-xxs12 and CaiT-xxs24, “fixed 1” that utilizes fixed overlap dimension 1 for all layers could significantly improve the accuracy by 2.39% and 3.58%, respectively. Furthermore, the accuracy is enormously enhanced by 7.41% for CaiT-xxs24 with “inc-1 (1)” which increases the overlap dimension by 1 every 1 layer with “1-indexed”. The effectiveness of our method is significant on CaiT models and we can ignore the slightly increased parameters and computations.
For Swin-Tiny, “fixed 1” paradigm could greatly enhance the accuracy by 1.23%, and the performance could be further boosted by a large margin with “dec-1 (1)” that decreases the overlap dimension by 1 every 1 layer with “1-indexed”. The effectiveness of overlapped heads is also demonstrated on the Transformer model with window-based hierarchical architecture.
Figure 5: The accuracy comparison during the training process. The accuracy comparison between the original method and our proposed approach for CaiT-xxs24 on val or test set of CIFAR-100, Tiny-ImageNet, and ImageNet is illustrated from left to right. The blue curve demonstrates our approach with the best performance and the red curve indicates the original method.
IV-D ImageNet
ImageNet-1k [11] includes 1k classes with more than 1 million training images and 50k validation images. By training and testing on ImageNet, we could investigate the effectiveness of our proposed approach on large datasets. The experimental results on ImageNet are demonstrated in Table VIII. “inc-0 (1)” and “inc-1 (1)” indicate the overlap dimension is increased by 1 every 1 layer with “0-indexed” and “1-indexed”, respectively. “dec-0 (1)” and “dec-1 (1)” represent the overlap dimension is decreased by 1 every 1 layer with “0-indexed” and “1-indexed”, respectively. Almost all models equipped with our proposed MOHSA have significant improvements on ImageNet.
TABLE VIII: The experimental results on ImageNet
Models Methods Accuracy Params FLOPs
ViT-Tiny original 69.25 5.7M 1.3G
fixed 1 69.11 5.8M 1.3G
fixed half 70.31 (+1.06) 6.2M 1.5G
inc-0 (1) 69.96 6.0M 1.4G
inc-1 (1) 69.70 6.1M 1.5G
dec-0 (1) 69.71 6.0M 1.4G
dec-1 (1) 69.34 6.1M 1.5G
ViT-Small original 77.24 22.0M 4.6G
fixed 1 77.57 22.1M 4.7G
fixed half 77.40 23.8M 5.3G
inc-0 (1) 77.87 22.6M 4.9G
inc-1 (1) 77.91 22.8M 4.9G
dec-0 (1) 78.03 (+0.79) 22.6M 4.9G
dec-1 (1) 77.99 22.8M 4.9G
CaiT-xxs12 original 67.56 6.6M 1.3G
fixed 1 67.61 6.6M 1.3G
fixed half 68.30 7.0M 1.6G
inc-0 (1) 68.63 6.7M 1.4G
inc-1 (1) 68.73 (+1.17) 6.7M 1.4G
dec-0 (1) 68.49 6.7M 1.4G
dec-1 (1) 68.25 6.7M 1.4G
CaiT-xxs24 original 70.56 11.9M 2.5G
fixed 1 73.64 12.0M 2.6G
fixed half 73.79 12.8M 3.1G
inc-0 (1) 74.26 (+3.70) 12.4M 2.8G
inc-1 (1) 74.19 12.4M 2.8G
dec-0 (1) 74.08 12.4M 2.8G
dec-1 (1) 74.17 12.4M 2.8G
Swin-Tiny original 78.52 28.3M 4.5G
fixed 1 78.44 28.4M 4.5G
fixed half 78.70 30.4M 5.0G
inc-0 (1) 78.70 29.4M 4.7G
inc-1 (1) 78.72 (+0.20) 29.6M 4.7G
dec-0 (1) 78.64 28.6M 4.7G
dec-1 (1) 78.61 28.8M 4.7G
ViT-Tiny could be boosted by more than 1% with “fixed half” that uses half of the head dimension as the overlap dimension for all layers. ViT-Small is enhanced by 0.79% with “dec-0 (1)” with negligible overhead. Additionally, utilizing 1 as the overlap dimension for all layers is not effective for ViT-Tiny on ImageNet. But using 1 as the overlap dimension is better than using half head dimension as the overlap dimension for all layers for ViT-Small.
Our approach has even more performance enhancement on CaiT models. CaiT-xxs12 is greatly improved by 1.17% with “inc-1 (1)” and the accuracy of CaiT-xxs24 is significantly increased by 3.70% with “inc-0 (1)”. For CaiT-xxs24, the accuracy is boosted by more than 3% by using only 1 as the overlap dimension for all layers, which illustrates the effectiveness and efficiency of our proposed method by remarkably boosting the performance of the models with minimum overhead.
TABLE IX: The ablation study on applying overlap to $Q$, $K$, $V$
Overlap ViT-Tiny CaiT-xxs12
Accuracy Params FLOPs Accuracy Params FLOPs
None 69.25 5.7M 1.3G 67.56 6.6M 1.3G
$Q$, $K$ 69.15 5.7M 1.3G 66.85 6.6M 1.3G
$V$ 69.62 6.0M 1.4G 68.53 6.7M 1.3G
$Q$, $K$, $V$ 69.96 6.0M 1.4G 68.63 6.7M 1.4G
For Swin-Tiny, simply using 1 as the overlap dimension is not effective on ImageNet. Even though increasing the overlap dimension to half of the head dimension improves the accuracy of Swin-Tiny, it achieves almost the same result as the paradigms of varying the overlap dimension by the depths of the layers which have fewer parameters and computations.
Table IX illustrates the ablation study on applying overlap paradigm “inc-0 (1)” to $Q$, $K$, or $V$ by using ViT-Tiny and CaiT-xxs12 on ImageNet. From the experimental results shown in Table IX, applying our approach to $V$ has more effect on boosting the performance, and applying the overlapping approach to $Q$, $K$, and $V$ could yield the best performance.
IV-E Analysis
In the experiments, we utilize different variants of our proposed MOHSA on various Vision Transformer models and datasets to illustrate the effectiveness of MOHSA. Overall, the enhancement of the models on various datasets is remarkable with such insignificant overhead. Our proposed approach might be exploited as a plug-and-play method for Vision Transformer models and might enhance the performance of those models with minimum extra cost. For different variants of our proposed method, the results manifest some differences. For a fixed paradigm that the overlap dimension is the same for all layers, the performance could be enhanced by only 1 overlap dimension in most cases and increasing the overlap dimension for fixed mode cannot guarantee better results. Additionally, the variants of varying overlap dimensions based on the depths of the layers demonstrate superior performance than the fixed paradigm on various models and datasets in most cases. Compared to the fixed paradigm with a high overlap dimension, varying overlap dimensions with the depths of the layers could save the number of parameters and computational costs.
To illustrate the accuracy during training, Fig. 5 manifests the accuracy of CaiT-xxs24 on the val or test set with the training proceeding between the original models and our proposed models. “inc-1 (3)”, “inc-1 (1)” and “inc-0 (1)” are selected as ours for comparison for CIFAR-100, Tiny-ImageNet, and ImageNet, respectively. The accuracy between the original models and our models demonstrates almost no difference at the early stage of training, while diverges significantly with the training proceeding.
Moreover, the best improvements for various datasets are mostly CaiT models. The possible reason might be that CaiT [15] utilizes talking-heads attention [30] before and after the softmax function. Equipped with our MOHSA, CaiT might have more effective information exchange between different heads and illustrate superior performance.
V Conclusion
In this paper, we have proposed a simple yet effective module MOHSA, to improve the original MHSA in Vision Transformers by overlapping the heads, allowing for information exchange during attention calculation in each head. Extensive evaluations and comparisons with the state-of-the-arts on multiple datasets have demonstrated the superior performance of the proposed approach. To the best of our knowledge, this is the first work to propose overlapping heads and achieve significant enhancement across various datasets for different Vision Transformer models. We hope our work will inspire the community to further explore the structure of Vision Transformers.